If 2x + 5= -25 and -3m - 6= 48, what is the product of x and m? (A). 270. (B). -33. (C). 3. (D). -270
Answer by rfer(16322)

2x+5=-25
2x=-30
x=-15
-----------
-3m-6=48
-3m=54
m=-18
----------
-15*-18=+270
answer (A)